Write an equation in slope-intercept form of each line described.
Slope 2; passes through (-5, 1)
Slope 1/3; passes through (6, 0)
Slope -2/5; passes through (5, 7)

Answer by Alan3354(69443)

Write an equation in slope-intercept form of each line described.
Slope 2; passes through (-5,1) *** don't put a space after the comma
---
Here's one way to do these.
Find b, the y-intercept using
y = mx + b
1 = 2*-5 + b
b = 11
--> y = 2x + 11
=====================================
Slope 1/3; passes through (6,0)
Use
y - y1 = m*(x - x1) (x1,y1) is the point (6,0)
y - 0 = (1/3)*(x - 6)
y = x/3 - 2
-------------
Slope -2/5; passes through (5,7)
Do the last one either way, or both ways to check.